# CLEAR-IR: Clarity-Enhanced Active Reconstruction of Infrared Imagery

### v1

@@ Abstract
This paper presents a novel approach for enabling robust robotic perception in dark environments using infrared (IR) stream. IR stream is less susceptible to noise than RGB in low-light conditions. However, it is dominated by active emitter patterns that hinder high-level tasks such as object detection, tracking and localisation. To address this, a U-Net-based architecture is proposed that reconstructs clean IR images from emitter-populated input, improving both image quality and downstream robotic performance. This approach outperforms existing enhancement techniques and enables reliable operation of vision-driven robotic systems across illumination conditions from well-lit to extreme low-light scenes.
@@ I Introduction
Conventional RGB cameras, while effective under well-lit conditions, suffer from significant degradation in low-light scenes [1]. Image quality is compromised due to sensor-level limitations, such as dark current and readout noise. Attempts to compensate using an increased exposure time introduce motion blur, rendering them ineffective for dynamic environments. Furthermore, data-driven approaches that require training networks in long-exposure or low-light image pairs add complexity and reduce generalisability across different camera models. In addition, such methods are fundamentally limited by information-theoretic constraints and if critical scene information is not captured by the sensor due to noise or extreme darkness, no amount of training data can fully reconstruct it. The effectiveness of these networks is therefore heavily dependent on both data availability and inherent scene observability.
@@
Another approach involves equipping robots with onboard illumination systems [2]. However, in certain rich environments such as underground tunnels or disaster zones, the Tyndall effect causes the light to scatter off airborne particles, saturating the camera view with undesirable visual artefacts and occluding key features in the scene. This method also suffers from overexposure, specular reflection, and increased power consumption. Moreover, the use of active illumination can produce strong shadows that obscure structural details and complicate downstream vision tasks such as object detection and localisation. To address these challenges, this work explores the use of infrared (IR) imaging, particularly active IR systems that operate independently of ambient illumination. As seen in Figure 1, the dark scene captured by the RGB camera fails to present good visual cues, whereas the same scene in the IR spectrum can be visualised with several objects in the scene. Although IR cameras are promising for dark-scene perception, they are not without their own limitations. Passive IR sensors, which operate without emitters, avoid structured-light artefacts but are typically expensive and limited to low-resolution outputs, making them impractical for robotics. Active IR systems, introduce structured light patterns that degrade image quality, interfere with feature extraction, and disrupt high-level perception tasks such as object detection, tracking, and SLAM.
To overcome this, a deep learning-based framework, CLEAR-IR (Clarity-Enhanced Active Reconstruction of Infrared Imagery), is proposed. The framework is designed to reconstruct clean and interpretable IR images from raw emitter-populated inputs. The method leverages a U-Net architecture trained to suppress structured light noise while preserving spatial and textural details critical for vision-based tasks. The output is a denoised, visually coherent IR image that can serve as a robust input to standard perception modules. This is the first work to leverage the infrared stream for high-level robotic tasks by removing active emitter patterns in scenes ranging from well-illuminated to extreme low-light. The contributions of this work are:
  • CLEAR-IR, a U-Net–based framework tailored for reconstructing clean IR images by suppressing structured-light artefacts, is introduced.
  • A composite loss function is developed to preserve spatial and textural details critical for robotic perception while robustly removing high-frequency emitter patterns.
  • Evaluation of IR image reconstruction for robotics is presented, demonstrating significant improvements across key tasks such as feature matching, object detection, and marker detection.
@@ II Related Work
Active infrared (IR) systems, which operate by illuminating a scene with an IR light source and capturing the reflected light, have been a viable alternative for dark-scene perception. These systems have been widely adopted in applications such as facial recognition [14] and depth sensing [15]. However, the structured light patterns they use for scene illumination and depth estimation, while beneficial for some tasks, present a significant challenge for 2D vision tasks like object detection and segmentation. Repetitive high-frequency patterns can be misinterpreted as texture or objects by standard vision algorithms, leading to false positives or missed detections, which limits the amount of work done using these sensors. Previous works have only attempted to remove fixed patterns from IR images, as seen in [16, 17, 18]. This just enhances the extremely noisy input of IR images, but does not focus on making it more useful for high-level robotics tasks.
The U-Net architecture, with its symmetric encoder-decoder structure and skip connections, has shown remarkable success in a variety of image-to-image translation tasks, and was first introduced for biomedical image segmentation in [19]. Its ability to capture both high-level semantic information from the encoder path and fine-grained spatial details from the decoder path makes it an ideal candidate for the specific denoising problem encountered in IR images. This architecture has also been successfully adapted for other image reconstruction tasks, such as super-resolution [20] and general image denoising [21].
In contrast to existing methods that primarily focus on low-light enhancement or general noise removal, this work addresses the specific and unique problem of structured light artefacts in active IR imagery. Using the U-Net architecture, it is possible to effectively learn to suppress these specific patterns while preserving the crucial underlying semantic information for high-level vision tasks.
@@ III Methodology
This section presents the CLEAR-IR framework. This method leverages a U-Net architecture and a composite loss function to effectively remove structured light artefacts while preserving critical visual details.
@@ III-A Emitter Pattern Characteristics
This variability in pattern appearance across different scenes creates a significant challenge in selecting adequate losses that can remove them from the images while being scene-agnostic.
III-B U-Net Architecture for Emitter Pattern Removal
To effectively reconstruct the image and remove the noise caused by the emitter pattern, a standard U-Net architecture was used. This convolutional neural network is particularly well suited for image-to-image translation tasks due to its symmetric encoder-decoder structure and strategic use of skip connections, as depicted in Figure 3.
The network is configured to process greyscale images with an input resolution of 640x480 pixels. The encoder path progressively downsamples the image, using a series of convolutional and max-pooling layers to extract increasingly abstract features and high-level contextual information.
The decoder path then reconstructs the image using transposed convolutional layers to upsample the features back to the original resolution. A crucial element of this architecture are its skip connections, which concatenate the feature maps from the encoder to the corresponding layers in the decoder. These connections are vital for preserving fine-grained spatial details that might otherwise be lost during the downsampling process. The network combines these features at each upsampling stage to produce a final, single-channel output image that is the same size as the input, with the emitter pattern effectively removed.
Figure 3: The U-Net architecture used for image reconstruction and emitter pattern removal.
III-C Dataset and Data Augmentation
A custom paired dataset of IR images and their corresponding clean ground-truth RGB images converted to greyscale was created for model training using an Intel Realsense D455 camera [22]. Data collection was performed in an indoor workspace with varied lighting conditions. The illumination levels and object distance were intentionally adjusted to capture emitter patterns with different sharpness and intensity, creating a varied dataset.
The initial dataset contained 10,000 paired images. To enhance model robustness and prevent overfitting, a data augmentation process was applied. For each image pair, five augmented versions were generated using random transformations, which included:
  • Rotation: Up to 30 degrees in either direction.
  • Flipping: Random horizontal and vertical flips.
  • Affine Transformations: Minor rotations and translations.
  • Zoom: Random zooming in by up to 30%.
  • Brightness and Contrast: Adjustments to simulate lighting variations.
The final data had 50000 paired sets of IR and greyscale RGB images and will be released as an open-sourced dataset.
III-D Composite Loss Function and Training
To produce high-quality image reconstructions, a composite loss function was used to guide the network’s training. This function is a weighted sum of multiple components, each designed to optimise a different aspect of the output, ensuring that the reconstructed image is not pixel accurate but perceptually and structurally similar to the ground truth. The main reason for choosing these losses besides mean absolute or mean squared error was due to the epipolar transformation of pixels between the IR and RGB images.
III-D1 Loss Components
The total loss is a weighted sum of several components. Let $\alpha,\beta,\gamma,\delta,\epsilon$, and $\zeta$ be the empirically tuned weighting coefficients for each corresponding loss component. The total loss is defined as:
[…]
  • Mean Absolute Error (MAE): This is a pixel-wise loss that minimises the average absolute difference between the predicted image $\hat{I}$ and the ground truth image $I$. It provides a foundational measure of pixel level accuracy.
    $\mathcal{L}_{mae}=\frac{1}{N}\sum_{i=1}^{N}|\hat{I}_{i}-I_{i}|$ (4)
  • Structural Similarity Index (SSIM) Loss: This loss measures the similarity in luminance, contrast, and structure between images. By minimising this loss, the network is encouraged to preserve the overall structural integrity of the scene [23].
    $\mathcal{L}_{ssim}=1-\text{SSIM}(\hat{I},I)$ (5)
  • High-Frequency and Sobel Edge Losses: These losses are used to preserve fine details and sharp boundaries. The high-frequency loss uses a Laplacian filter to penalise differences in the high-frequency components of the images, while the Sobel edge loss penalises differences in image gradients to ensure edges are accurately reproduced.
    $\mathcal{L}_{freq}=\left\|(\hat{I}-I)*K_{Laplacian}\right\|_{1}$ (6)
    $\mathcal{L}_{sobel}=\left\|\nabla_{x}\hat{I}-\nabla_{x}I\right\|_{1}+\left\|\nabla_{y}\hat{I}-\nabla_{y}I\right\|_{1}$ (7)
  • Perceptual Loss: This loss component uses a pre-trained VGG19 network as a feature extractor. The network’s weights, which were pre-trained on the ImageNet dataset [24], are frozen during training. This allows the model to leverage learned feature representations that are more aligned with human visual perception than simple pixel values. To apply this to greyscale images, the input is converted to a 3 channel format by concatenating the single channel three times. Instead of comparing pixels directly, this loss compares the feature maps from a set of specific intermediate layers within the VGG19 network. By using layers such as block3_conv2, block4_conv2, and the layers within block5, the loss encourages the reconstructed image to have similar high level content and style as the ground truth. This is crucial for generating an output that is visually convincing and suitable for high level vision tasks [25].
    $\mathcal{L}_{perceptual}=\sum_{j}\|\Phi_{j}(\hat{I})-\Phi_{j}(I)\|_{1}$ (8)
    where $\Phi_{j}(\cdot)$ is the feature map from layer $j$ of the VGG19 network.
  • Total Variation (TV) Loss: This regularizer promotes spatial smoothness and helps to reduce spurious noise in the output image [26].
    $\mathcal{L}_{tv}=\sum_{i,j}\left(\left(\hat{I}_{i+1,j}-\hat{I}_{i,j}\right)^{2}+\left(\hat{I}_{i,j+1}-\hat{I}_{i,j}\right)^{2}\right)$ (9)
III-D2 Weighting Strategy
The weighting coefficients for each component ($\alpha,\beta,\gamma,\delta,\epsilon,\zeta$) were experimentally tuned to balance their influence on the final output. The primary goal was to prioritise visual quality and structural integrity over a strict pixel-by-pixel match. High weights were assigned to $\mathcal{L}_{ssim}$ to preserve scene structure, while a significant but not overwhelming weight was given to $\mathcal{L}_{perceptual}$ to ensure the output was visually convincing. Moderate weights were used for $\mathcal{L}_{freq}$ and $\mathcal{L}_{sobel}$ to preserve fine details and edges while also not prioritising emitter patterns which can be classed as high-frequency patterns in the image. In contrast, very low weights were given to $\mathcal{L}_{mae}$ and $\mathcal{L}_{tv}$, as their purpose was to act as subtle regularizers rather than primary training objectives.
III-E Preprocessing and Training
After augmentation, the dataset, which contained 50,000 paired images, 90% of the data was split into a training set and the remaining 10% into the validation set. All images were resized to a consistent resolution of 640x480 pixels and normalised to the range of [0, 1].
The model was trained using the Adam optimizer with a CosineDecayRestarts learning rate schedule to minimise the composite loss function. To prevent overfitting, an early stopping mechanism with a patience of 30 epochs was implemented.
The model was trained for a total of 126 epochs, at which point it met the early stopping criteria, indicating that a sufficient level of convergence had been reached.
IV Experimental Results
This section presents a comprehensive evaluation of the proposed CLEAR-IR framework, focusing on its effectiveness in preparing infrared images for high-level computer vision tasks. Since this is the first work to specifically address the removal of active emitter patterns, the primary evaluation is a qualitative and quantitative comparison of the performance of enhancement techniques on the RGB stream versus the denoised images of the IR stream produced by this work.
IV-A Qualitative and Quantitative Analysis
To provide a visual assessment of the method’s reconstruction quality and its robustness in low light conditions, a qualitative comparison is first presented of a raw IR input, the model’s reconstructed output, and the corresponding ground truth image, as seen in Figure 4. The method’s output is also compared against state-of-the-art low-light RGB image enhancement techniques to highlight the inherent advantages of using the IR spectrum in these conditions.
(a) Raw IR Input
(b) Denoised Output
(c) Ground-Truth
Figure 4: A qualitative comparison of the raw IR input with emitter patterns, the reconstructed image, and the corresponding ground-truth image.
As seen in the visual examples, the reconstructed output successfully removes the emitter patterns while preserving key features such as object boundaries and textures. The output image is visually coherent and appears cleaner, making it a much more suitable input for downstream computer vision algorithms. Even object shadows which are hard to see in the raw IR image can be seen clearly in the reconstructed image.
The advantages of the CLEAR-IR method are further emphasised through a quantitative comparison with traditional methods such as Contrast Limited Adaptive Histogram Equalisation (CLAHE) [27], Retinex Theory inspired techniques [28] and state-of-the-art low-light RGB image enhancement techniques such as Zero-Reference Deep Curve Estimation (Zero-DCE) [29], Color and Intensity Decoupling Network (CIDNet) [30], and LLFormer [31]. For this comparison, the similarity of each method’s output to a ground truth RGB image is evaluated using the Structural Similarity Index Measure (SSIM).
The visual results in Figure 5 shows the outputs from different techniques on the scene. While other methods attempt to enhance the image, the enhanced RGB images are affected by the effects of noise observed in low-light scenes. The IR cameras are more robust to the effect of this noise in low-light, as a result, the CLEAR-IR framework produces a clean, artifact-free image that maintains the scene’s structural integrity, as reflected by its superior SSIM score. This demonstrates that the proposed method is more effective at generating a representation of the scene suitable for downstream vision tasks.
TABLE I: Quantitative comparison of image enhancement methods against ground truth.
Method SSIM
Retinex Theory [27] 0.7418
CLAHE [28] 0.5672
Zero-DCE [29] 0.7624
LLFormer [30] 0.7063
CIDNet [31] 0.7788
CLEAR-IR (Proposed Method) 0.9521
The results, presented in Table I, show that the denoised IR output achieves a significantly higher SSIM score and is capable of maintaining the true structural integrity of the scene, even when compared to methods that produce a colour output.
Beyond the general quality metrics, the true utility of the proposed CLEAR-IR framework is demonstrated by its ability to facilitate high-level robotics tasks that are otherwise compromised by the structured light patterns and noise inherent in raw IR images and RGB images respectively. Table II presents a comprehensive quantitative comparison of the performance of various image enhancement methods on three high level tasks such as ORB feature matching, object detection, and ArUco marker detection. The results highlight a critical distinction between methods that simply improve image aesthetics and those that produce a robust, clean representation suitable for downstream algorithms.
@@ IV-A Qualitative and Quantitative Analysis
(h) CLEAR-IR (Proposed Method)
Figure 5: Visual comparison of a ground truth RGB scene, a dark RGB scene, and the outputs of various image enhancement methods, including the proposed Denoised IR method.

### v2

@@ Abstract
This paper presents a novel approach for enabling robust robotic perception in dark environments using infrared (IR) stream. IR stream is less susceptible to noise than RGB in low-light conditions. However, it is dominated by active emitter patterns that hinder high-level tasks such as object detection, tracking and localisation. To address this, a Deep Multi-scale Aware Overcomplete (DeepMAO) inspired architecture is proposed that reconstructs clean IR images from emitter populated input, improving both image quality and downstream robotic performance. This approach outperforms existing enhancement techniques and enables reliable operation of vision driven robotic systems across illumination conditions from well-lit to extreme low-light scenes. The results outline the ability of this work to be able to mimic RGB styling from the scene and its applicability on robotics tasks that were trained on RGB images, opening the possibility of doing these tasks in extreme low-light without on-board lighting.
@@ I Introduction
Conventional RGB cameras, while effective under well-lit conditions, suffer from significant degradation in low-light scenes [14]. Image quality is compromised due to sensor-level limitations, such as dark current and readout noise. Attempts to compensate using an increased exposure time introduce motion blur, rendering them ineffective for dynamic environments. Furthermore, data-driven approaches relying solely on visible spectrum low light and long exposure pairs are fundamentally limited by information-theoretic constraints. If critical scene information is not captured by the sensor due to extreme darkness or sensor level noise, it cannot be reconstructed regardless of the training data volume. Consequently, the efficacy of these networks remains strictly bound by the inherent observability of the visible spectrum in degraded conditions.
@@
Another approach involves equipping robots with onboard illumination systems [6]. However, in certain rich environments such as underground tunnels or disaster zones, the Tyndall effect causes the light to scatter off airborne particles, saturating the camera view with undesirable visual artefacts and occluding key features in the scene. This method also suffers from overexposure, specular reflection, and increased power consumption. Moreover, the use of active illumination can produce strong shadows that obscure structural details and complicate downstream vision tasks such as object detection and localisation. To address these challenges, this work explores the use of infrared (IR) imaging, particularly active IR systems that operate independently of ambient illumination. As seen in Figure 1, the dark scene captured by the RGB camera fails to present good visual cues, whereas the same scene in the IR spectrum can be visualised with several objects in the scene. Although IR cameras are less susceptible to such scattering due to its longer wavelength and are promising for dark-scene perception, they are not without their own limitations. Passive IR sensors, which operate without emitters, avoid structured-light artefacts but are typically expensive and limited to low-resolution outputs. Conversely, active IR systems often rely on structured light or flash patterns, and while these provide better penetration through dusty scenes than having an ambient light in extreme low-light scenes, the resulting high-intensity reflections and projected patterns can interfere with feature extraction, and disrupt downstream perception tasks such as object detection and Simultaneous Localisation and Mapping (SLAM).
@@
To overcome this, a deep learning-based framework, CLEAR-IR (Clarity-Enhanced Active Reconstruction of Infrared Imagery), is proposed. The framework is designed to reconstruct clean and interpretable IR images from raw emitter populated inputs. The method leverages a Deep Multi-scale Aware Overcomplete (DeepMAO [26]) inspired architecture trained to suppress structured light noise while preserving spatial and textural details critical for vision-based tasks. The output is a reconstructed, visually coherent IR image that can serve as a robust input to standard perception modules. This is the first work to leverage the infrared stream for high-level robotic tasks by removing active emitter patterns in scenes ranging from well-illuminated to extreme low-light. The contributions of this work are:
  • CLEAR-IR, a DeepMAO inspired architecture tailored for reconstructing clean IR images by suppressing structured-light artefacts, is introduced.
  • We utilise a composite loss function specifically selected to accommodate the spectral disparity and lack of pixel perfect alignment between IR and RGB inputs, allowing for robust pattern removal while preserving details critical for robotic perception.
  • Evaluation of IR image reconstruction for robotics is presented, demonstrating significant improvements across key tasks such as object detection and marker detection.
  • Comparison with other SOTA techniques for localisation in extreme low-light scenes.
@@ II Related Work
Active infrared (IR) systems provide a robust alternative for dark-scene perception by illuminating the environment with a dedicated source. While high power IR floodlights offer uniform coverage, they are often impractical for mobile robotics due to the inverse-square law, the power required for an adequate signal-to-noise ratio scales quadratically with distance. Furthermore, floodlights frequently cause shadows and also severe near-field saturation, where excessive photon flux annihilates local contrast and occludes critical features. To mitigate this while providing the structural texture necessary for navigation on featureless surfaces, depth camera manufacturers utilise structured-light or modulated laser patterns [1]. Beyond their sensing advantages, these active IR systems are already standard components in many robotic platforms and leveraging them for 2D perception minimizes system complexity, reduces payload weight, and lowers maintenance costs compared to adding auxiliary illumination hardware. By distributing energy into discrete, evenly spaced points, these systems maintain a higher usable dynamic range even as pattern density increases at close proximity, the localised nature of the emission prevents the global washout characteristic of floodlights. However, these high-frequency patterns present a significant challenge for 2D vision tasks, as standard algorithms often misinterpret them as scene texture, leading to false detections. Previous denoising works [12, 21, 2] focus primarily on removing static sensor level artefacts rather than addressing the dynamic, pattern dependent noise inherent to these active emitters. Simply enhancing these inputs does not bridge the domain gap required for robotics, as static corrections cannot account for the characteristics discussed in Section III-A.
The U-Net architecture, with its symmetric encoder-decoder structure, skip connections, has shown remarkable success in a variety of image-to-image translation tasks, and was first introduced for biomedical image segmentation in [24]. Its ability to capture both high-level semantic information from the encoder path and fine-grained spatial details from the decoder path makes it an ideal candidate for the specific reconstruction problem encountered in IR images. This architecture has also been successfully adapted for other image reconstruction tasks, such as super-resolution [16] and general image denoising [34].
In contrast to existing methods that primarily focus on low-light enhancement or general noise removal, this work addresses the specific and unique problem of structured light artefacts in active IR imagery. Using the DeepMAO architecture, it is possible to effectively learn to suppress these specific patterns while preserving the crucial underlying semantic information for high-level vision tasks.
@@ III Methodology
This section presents the CLEAR-IR framework. Unlike standard encoder-decoder approaches, CLEAR-IR uses a hybrid dual-branch architecture to effectively remove structured light artefacts while preserving critical visual details.
@@ III-A Emitter Pattern Characteristics
This variability in pattern appearance across different scenes poses a challenge in selecting adequate losses that can remove them from the images while being scene agnostic.
III-B CLEAR-IR Architecture
To preserve high-frequency spatial details, this work adopts a modified DeepMAO architecture. Although the original DeepMAO was modelled for satellite image segmentation using a pretrained EfficientNet backbpne along with a combination of focal loss and dice loss, this work adapts its dual-stream ideology for IR image reconstruction. This work simplifies the original architecture by replacing the heavy encoder with a U-Net backbone and we simplify the multi-scale overcomplete (MSO) module from non-integer interpolations to a sequence of full-resolution convolution layers. We retain the original architecture’s additive residual fusion to facilitate a direct gradient flow for fine-grained textures. Furthermore, we transition the training objective from pixel-level classification to intensity regression, utilising a combined perceptual and structure loss tailored for scene reconstruction.
III-B1 Context Stream (U-Net Backbone)
Unlike the original DeepMAO which relied on a heavy EfficientNet-B3 encoder, our first branch utilises a U-Net backbone designed to efficiently capture global context and low-frequency structural information.
  • Hierarchical Encoding: The encoder processes the input through a series of convolutional blocks with increasing filter depths (32, 64, 128) and $2\times 2$ max-pooling layers. This effectively expands the receptive field, allowing the network to understand global scene geometry.
  • Regularisation: To ensure stability during training and prevent overfitting on specific emitter patterns, dropout with a rate of $0.2$ is integrated into the deeper layers of the encoder and bottleneck.
  • Decoder and Skip Connections: The decoder utilises transposed convolutions to upsample feature maps, concatenated with corresponding features from the encoder to recover spatial information lost during pooling.
@@ III-B1 Context Stream (U-Net Backbone)
(h) Raw IR
(i) U-Net
(j) CLEAR-IR (Proposed Method)
Figure 3: Visual comparison of enhancement methods on example scene.
III-B2 Detail Stream (Overcomplete Branch)
While the pooling operations in the U-Net are effective at smoothing out the artificial structured light patterns, they are also destructive to the true high frequency details of the scene. To mitigate this, the second branch acts as an overcomplete pathway.
  • Full Spatial Resolution: In the original DeepMAO, the overcomplete path utilized complex non-integer interpolations and dilated convolutions. We simplify this mechanism to better suit image reconstruction. Our branch accepts the initial feature maps (32 filters) and processes them through a cascade of convolutional layers without any scaling, pooling, or strided operations.
  • Detail Preservation and Noise Suppression: By strictly maintaining a 1:1 input-to-output spatial correspondence, this branch bypasses the destructive downsampling of the U-Net. It specialises in isolating the true high-frequency structural boundaries from the structured light noise, ensuring that sharp edges and fine scene textures are preserved and injected back into the final reconstruction.
III-B3 Feature Fusion
The outputs of the U-Net branch and the spatial overcomplete branch are integrated using an additive fusion strategy, mirroring the residual nature of the original network.
$F_{fused}=F_{unet}\oplus F_{overcomplete}$ (1)
where $\oplus$ denotes element-wise addition. This residual-style fusion allows the network to combine the clean structural predictions from the U-Net with the texture corrections from the overcomplete branch. The fused features are passed through a final convolutional refinement block to generate the continuous intensity values of the reconstructed IR image.
III-C Loss Function
Despite the extrinsic calibration and warping process, minor alignment residuals persist due to parallax and sensor baseline between the IR and RGB camera. Therefore, this work uses a composite loss function that prioritises perceptual quality and structural integrity over simple pixel-wise accuracy. The total loss $\mathcal{L}_{\text{total}}$ is a weighted sum of six components:
[…]
The components are defined as follows:
  • Spatial Fidelity ($\mathcal{L}_{mae},\mathcal{L}_{ssim}$): This work use Mean Absolute Error (MAE) for base intensity matching and Structural Similarity Index (SSIM) [30] to preserve luminance and contrast structure.
  • Detail Preservation ($\mathcal{L}_{freq},\mathcal{L}_{sobel}$): To retain sharp edges, $\mathcal{L}_{freq}$ minimizes the $L_{1}$ distance between Laplacian-filtered images, while $\mathcal{L}_{sobel}$ penalizes discrepancies in image gradients.
  • Perceptual Quality ($\mathcal{L}_{perceptual}$): A VGG19 network pre-trained on ImageNet [7] is used as a feature extractor. The loss minimises the $L_{1}$ distance between feature maps of specific intermediate layers, ensuring the reconstruction aligns with human visual perception [18]:
  • Regularization ($\mathcal{L}_{tv}$): Total Variation (TV) loss promotes spatial smoothness to reduce spurious noise in uniform regions.
The weighting coefficients ($\alpha,\beta,\gamma,\delta,\epsilon,\zeta$) were experimentally tuned to prioritize $\mathcal{L}_{ssim}$ and $\mathcal{L}_{perceptual}$, ensuring the model focuses on structural removal of the pattern rather than overfitting to minor pixel misalignments.
III-D Data Augmentation and Training Implementation
A custom dataset was developed by capturing 6,719 initial pairs of active IR and greyscale RGB images using an Intel RealSense D455 camera [17]. To ensure spatial correspondence, we utilised factory calibrated extrinsics to align the sensors. These base frames were diversified through data augmentation, including random rotations ($\pm 30^{\circ}$), flipping, affine transformations, and brightness adjustments resulting in a final dataset of 33,595 images. The model was trained to map IR inputs to an RGB-style ground truth, enabling the deployment of downstream vision tasks originally developed for the visible spectrum. The network was trained on a 90 to 10 data split at $640\times 480$ resolution using the Adam optimizer with a cosine decay scheduler. Early stopping was also used to ensure convergence while preventing overfitting.
IV Experimental Results
This section presents a comprehensive evaluation of the proposed CLEAR-IR framework, focusing on its effectiveness in preparing infrared images for high-level computer vision tasks. Since this is the first work to specifically address the removal of active emitter patterns, the primary evaluation is a qualitative and quantitative comparison of the performance of enhancement techniques on the RGB stream versus the reconstructed images of the IR stream produced by this work.
The advantages of the CLEAR-IR method is emphasised through a quantitative comparison against both traditional baselines and a diverse set of state of the art (SOTA) enhancement methods. We evaluate against classical spatial-domain techniques, including CLAHE [27] and retinex theory inspired methods [3]. To benchmark against modern deep learning advances, this work evaluates Zero-DCE [13], a zero-reference method for unsupervised curve estimation, LLFormer [29], a benchmark Transformer-based architecture optimised for long-range dependency modelling and CIDNet [31], a multi-axis decoupling network that utilises a new HVI colour space to suppress colour bias and noise. These methods provide a robust evaluation across the leading unsupervised, attention-based, and colour-decoupling methodologies.
The visual results in Figure 3 shows the outputs from different techniques on an unknown scene. While other methods attempt to enhance the image, the enhanced RGB images are affected by the effects of noise observed in low-light scenes. The IR cameras are more robust to the effect of this noise in low-light, as a result, the CLEAR-IR framework produces a clean, artifact free image that maintains the scene’s structural integrity without introducing any additional artefacts or hallucinating. This demonstrates that the proposed method is more effective at generating a representation of the scene suitable for downstream vision tasks.